OpenStudy (anonymous):

solve each system by graphing y = x^2 + 1 y = x + 1 can someone help me with these??

OpenStudy (anonymous):

@jim_thompson5910 can you please help me with graphing?

jimthompson5910 (jim_thompson5910):

do you know how to graph each equation?

OpenStudy (anonymous):

no i dont actually

jimthompson5910 (jim_thompson5910):

what kind of calculator do you have?

OpenStudy (anonymous):

just the one on my computer i guess

jimthompson5910 (jim_thompson5910):

ok, I recommend downloading and installing geogebra (it's free)

jimthompson5910 (jim_thompson5910):

that's the main graphing program I use

jimthompson5910 (jim_thompson5910):

let me know once you have that set up

OpenStudy (anonymous):

The solution to the system of linear equations will be where they intersect (that is, where they share the same x and y value).

OpenStudy (anonymous):

If you're using a Mac, use Grapher. It's simple. See attached.

OpenStudy (anonymous):

To open Grapher, hit Command-Space, and type "Grapher," then press Enter. Select 2D Graph. Then, Go to the Equation menu and insert your equations.

OpenStudy (anonymous):

comand space?

OpenStudy (anonymous):

The little command button by the space bar. It looks like a square with loops at the corners. Commad-space opens the search function on your mac; press them together and the top right search box should open. Type Grapher in it.

OpenStudy (anonymous):

See attached for the search box (spotlight)

OpenStudy (anonymous):

Alternatively, you could pinch three fingers and your thumb together on your trackpad to open your applications, go to utilities, and select grapher: see attached

OpenStudy (anonymous):

but it didnt show up? @digitallogic

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

its probably blocked since this is a school computer

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

is there any other way do this have this on apps?

OpenStudy (anonymous):

like on a phone?

OpenStudy (anonymous):

What kind of phone do you have?

OpenStudy (anonymous):

android galaxy s4

OpenStudy (anonymous):

and an iphone 4

OpenStudy (anonymous):

Looks like it has good reviews.

OpenStudy (anonymous):

im sorry but i cant open it, whats it called?

OpenStudy (anonymous):

Oh lol. This android one was called Graphing Calculator by Mathlab

OpenStudy (anonymous):

okay thanks so much!

OpenStudy (anonymous):

The apple one is called Free Graphinc Calculator by William Jockusch

OpenStudy (anonymous):

Graphing*

OpenStudy (anonymous):

Alright good luck :)

OpenStudy (anonymous):

Here:

OpenStudy (anonymous):

So another way to solve this system without graphing is to just set them equal to each other and solve for x. You are essentially finding what values of x make both equations equal to each other.

OpenStudy (anonymous):

oh thanks i have to solve them with graphing though, but when i typed in y=x^2+1 and y=x+1 they just showed me a line going straight though the middle diagonally

OpenStudy (anonymous):

Look at the picture I attached two posts up.

OpenStudy (anonymous):

yeah i saw it but i wanted to try and match it to see if i was doing it correct but thanks

OpenStudy (anonymous):



OpenStudy (anonymous):

that's a little off but it should look that. when x=0 and x=1 the lines intersect

OpenStudy (anonymous):

so your points are (0, y) and (1,y_2)

OpenStudy (anonymous):

you found that from the graph; now plug in 0 into the easy equation (x+1) and you get y=1, plug in 1 and you get y=2

OpenStudy (anonymous):

Yeah that picture from Grapher on my computer is just squished a little. It's the same graph, just a different scale.

OpenStudy (anonymous):

I could set the scale to math the graph from wolfram and it would be exactly the same; both are correct

OpenStudy (anonymous):

so your answers are (0,1) and (1,2)

OpenStudy (anonymous):

those are the points of intersection; the solutions to the system are x=0, x=1

OpenStudy (anonymous):

which you found from the graph

OpenStudy (anonymous):

where are you? europe?

OpenStudy (anonymous):

it's 22:42 here lol; you're in school, i am in a coffee shop

OpenStudy (anonymous):

no im in america haha no im at home but its easier to take online classes

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

its 11;41 pm here

OpenStudy (anonymous):

22;42?

OpenStudy (anonymous):

i said it that way cause i assumed you are european; they use 24 hour time, mostly

OpenStudy (anonymous):

22;42 = 10:42PM

OpenStudy (anonymous):

10PM is th 22nd hour in the day; 24 hours in a day, midnight is 00:00, 11:59 pm is 23:59

OpenStudy (anonymous):

i wish i was in europe lol

OpenStudy (anonymous):

no AM/PM nonsense lol

OpenStudy (anonymous):

just the same, they use actual mass for measuring "weight"

OpenStudy (anonymous):

ive lived my whole life using am and pm and yeah i knew about that haha are you american also?

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

i studied abroad in europe for a year tho

OpenStudy (anonymous):

i wish i could go there and study

OpenStudy (anonymous):

someday! do well in school and you will have the opportunity to study abroad in college for sure

OpenStudy (anonymous):

even if it costs something, but it is possible

OpenStudy (anonymous):

im always leaving to the middle east every year so money isnt a big problem im just not that smart in school tbh

OpenStudy (anonymous):

oh cool :)

OpenStudy (anonymous):

oh yeah you are, it just takes practice and effort

OpenStudy (anonymous):

a hard-working b student will be more successful than a lazy genius

OpenStudy (anonymous):

no matter what

OpenStudy (anonymous):

and a hard-working b student will become an a student through practice

OpenStudy (anonymous):

i wish i was a b student right now i have a c in algebra because i cant get pass this lesson of graphing lol

OpenStudy (anonymous):

lol well hey, i made a c in algebra back in the day